### v1

@@ 3.1 Formulation and Architectural Overview
The ground-based cloud image sequence extrapolation task is formulated as follows: given an input sequence of $T$ historical frames, denoted as $X_{t}^{T}=\left\{x_{i}\right\}_{t+1}^{T}$, where each frame $x_{i}\in\mathbb{R}^{C\times H\times W}$, the objective is to predict a sequence of $\tau$ future frames after $T$, $Y_{T+1}^{T+\tau}=\left\{y_{i}\right\}_{T+1}^{T+\tau}$, where $y_{i}\in\mathbb{R}^{C\times H\times W}$. The model, parameterized by $\theta$, learns a mapping $F_{\theta}:X_{t}^{T}\rightarrow Y_{T+1}^{T+\tau}$ by maximizing the log-likelihood of the predicted frames (e.g., cloud images) with respect to the ground-truth counterparts.
Specifically, $C$, $H$, and $W$ correspond to the channel, height, and width dimensions of the images, respectively. Notably, the input and output of cloud extrapolation tasks are structured as tensors, i.e., $X_{t}^{T}\in\mathbb{R}^{T\times C\times H\times W}$ and $Y_{T+1}^{T+\tau}\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the frame rate. Formally, our objective prediction optimization model can be parameterized as $\theta$:
$\theta_{T}=arg\underset{\theta}{max}{\textstyle\sum_{i=T+1}^{T+\tau}}logP\left(y_{i}\mid x_{i};\theta\right)$ (1)
where the predictive model constitutes a learnable mapping $\theta_{T}$ that maximizes the log-likelihood between predicted cloud images and the ground-truth counterparts.
@@ 3.2 Overview of Structure
The proposed USF-Net, as shown in Fig. 2 (a), adopts a universal encoder-decoder architecture. The procedure of our USF-Net is described in Algorithm 1. A unified spatiotemporal module is introduced at the bottleneck to enhance the accuracy of ground-based cloud image sequence extrapolation by explicitly incorporating temporal guidance. As shown in Fig. 2 (a) and (b), in the encoder part, the input $X_{t}^{T}\in\mathbb{R}^{B\times T\times C\times H\times W}$ is first processed by three basic layers. Each layer consists of a $3\times 3$ DW convolutional layer for local feature extraction, followed by layer normalization, a DW convolutional layer, a Squeeze-and-Excitation (SE) block and a convolutional layer. The residual connection is used to enrich the representation capabilities. We denote the feature map of each layer as $f_{i}(i\in[1,2,3])$. The size of $f_{i}$ is $T\times 2^{i-1}\cdot 64\cdot\left(H\times W\right)/2^{i+1}(i\in[1,2,3])$, ultimately generating an intermediate feature map $X_{B}\in T\cdot C_{3}\cdot\left(H\times W\right)/16$. This feature map $X_{B}$ is then fed into the USTM, which comprises dual spatial and temporal branches coupled with a DSM. The spatial branch is employed to extract multi-scale spatial information, while the temporal branch is employed to capture temporal dependencies from the input sequence. These temporal cues are subsequently channeled through the DSM to guide spatial feature fusion, achieving unified spatiotemporal modeling. In the decoder part, to mitigate the “ghosting effect” caused by contextual information loss, a novel gated unit termed the DUM is designed to enable temporally guided spatial feature refinement. Within DUM, initial temporal features are reweighted via a gate mechanism to refine multi-scale spatiotemporal representations, preserving contextual coherence. It is noteworthy that USF-Net jointly extracts spatial features and temporal information, leveraging temporal flows to guide the generation of high-precision semantic representations. Consequently, our architecture can be interpreted as a spatiotemporally unified framework optimized for ground-based cloud image extrapolation, balancing computational efficiency with prediction accuracy. The details of our method are as follows.
@@
3.3 Unified SpatioTemporal Module
Learning multi-scale contextual information from scale-variant cloud imagery is critical for ground-based cloud image extrapolation tasks, as it facilitates precise spatial feature extraction. In real-world scenarios, complex meteorological conditions, such as variable wind speeds and diverse atmospheric flow patterns, result in non-stationary and nonlinear motion dynamics within cloud sequences. Therefore, establishing effective spatiotemporal dependencies is essential to achieve accurate predictions in cloud image extrapolation tasks.
However, most existing cloud image extrapolation methods that apply large-scale kernels ignore the importance of dynamic adaptive selection in capturing multi-scale fine-grained features from cloud imagery. In addition, spatiotemporal feature extraction in these approaches is often decoupled, lacking interactive guidance between spatial and temporal representations, and ignoring the influence of temporal information flows on spatial feature learning. Furthermore, prevalent methods employing self-attention mechanisms or LSTM-based recurrent architectures for temporal modeling incur prohibitive computational complexity due to their quadratic or sequential operational paradigms, limiting practical applicability in resource-constrained scenarios.
To this end, we design a unified spatiotemporal module to improve the ability to explore multi-scale contextual information of cloud imagery adaptively. Temporal dependencies are efficiently captured using a low-complexity attention mechanism. By incorporating TGM to inject temporal flows into spatial representations, the model ensures holistic spatiotemporal dependency modeling, significantly improving the precision of ground-based remote sensing cloud image extrapolation.
[…]
3.3.1 Spatial Information Branch
Cloud formations in natural meteorological conditions often exhibit complex nonlinear and non-stationary motion, leading to multi-scale variations in cloud image extrapolation tasks. To achieve precise predictions for ground-based remote sensing cloud image extrapolation, comprehensive and dynamically adaptive multi-scale contextual information is essential.
To this end, a spatial information branch with SSM is proposed to enhance the capability of dynamically extracting multi-scale contextual features, as shown in Fig. 3. Inspired by the Large Kernel Selection (LSK) mechanism [LSK], the dynamic adaptive large convolutional kernel adaptively adjusts the receptive field for each spatial target based on the scale of cloud formations. Specifically, the SSM is passing through a DW convolutional layer and a GeLU activation function to balance feature extraction efficiency with nonlinear representation capacity. To enhance the ability of the network to focus on the most relevant spatial contextual regions for multi-scale cloud formation extraction, multi-scale convolutional kernels are employed to adaptively select spatial features, where each kernel $K_{i}$ generates a corresponding feature map $X_{K_{i}}$. This process can be formalized as:
$X_{K_{i}}=DW\left(K_{i}\right)$ (2)
where $i$ denotes the number of large convolutional kernels. In this paper, kernels with configurations $K_{1}=3$, $d_{1}=1$ (dilation rate) and $K_{2}=7$, $d_{2}=3$ are utilized. As demonstrated by the following formula:
$RF_{1}=k_{1}$ (3)
$RF_{2}=d_{2}\left(k_{2}-1\right)+RF_{1}$ (4)
the proposed selection module effectively performs an explicit decomposition of a large kernel ($K=21$). By progressively increasing the kernel size and dilation rate, the explicitly decomposed convolution operations generate varying receptive field sizes, enhancing the network’s multi-scale representational capacity while significantly reducing parameter count.
@@ 3.3.1 Spatial Information Branch
Subsequently, the multi-scale feature maps are fused to generate $X_{K}$, which encapsulates diverse receptive fields. To enable interaction between cross-spatial features, channel-wise average pooling and max pooling are applied to $X_{K}$, followed by a convolutional layer (Conv) to produce the spatial interaction attention map $\widetilde{SA}$. This process is mathematically formalized as:
$X_{K}=Cat\left(X_{K_{1}},X_{K_{2}}\right)$ (5)
$\widetilde{SA}=Conv\left(Avg\left(X_{K}\right),Max\left(X_{K}\right)\right)$ (6)
@@
where $Cat$ denotes the concatenation of the channel, $Avg$ and $Max$ denote the average pooling and max pooling, respectively.
To achieve dynamic adaptive extraction of multi-scale cloud formations, the attention operators $\widehat{SA}_{i},i\in\left[1,2\right]$ for distinct receptive fields are derived by applying a sigmoid function to $\widetilde{SA}$. The feature maps from the decomposed large-kernel sequence are weighted by their corresponding selection weights, and the fused features are processed via a Conv layer to obtain the multi-scale spatial selection attention map $X_{SA}$. Finally, the spatial branch’s output $X_{S}$ is computed by combining $X_{B}$ with $X_{SA}$, as expressed in:
@@
$X_{SA}=Conv\left(Cat\left(\widehat{SA}_{1}\ast X_{K_{1}}\widehat{SA}_{2}\ast X_{K_{2}}\right)\right)$ (8)
$X_{S}=X_{SA}\otimes X_{B}$ (9)
where $\otimes$ denotes the element-wise production. By dynamically adjusting the receptive fields of spatial targets within the spatial branch, the proposed method effectively captures contextual information across varying cloud scales.
3.3.2 Temporal Information Branch
Cloud exhibits non-stationary motion over time due to meteorological and temporal influences, necessitating the acquisition of temporal information from cloud imagery sequences to model their motion trends accurately. Conventional approaches often employ recurrent networks or attention mechanisms to extract temporal dependencies from sequences. However, recurrent networks suffer from an inability to parallelize data processing, impotent the rendering of real-time, precise prediction due to the prohibitive computational costs. Self-attention mechanisms in most existing methods introduce quadratic complexity, resulting in substantial computational overhead. Therefore, inspired by agent attention, we propose a TAM, which synergizes the high precision of Softmax attention and the low complexity of Linear attention, achieving a favorable trade-off between computational efficiency and representational capacity for spatiotemporal modeling.
As shown in Fig. 4 (a), the proposed temporal information branch comprises two stacked components: conv embedding (CE) and TAMs. Different from conventional Vision Transformers (ViTs) that partition images into non-overlapping patches via linear projections, we select a CNN layer that replaces the patch embedding (PE) layer, as this design choice mitigates spatial information degradation and preserves multi-scale contextual coherence in cloud imagery. This method inherently retains spatial and positional information without requiring auxiliary positional encoding. The input channel dimension aligns with the embedding dimension in ViT architectures. Specifically, as shown in Fig. 4 (b), a CBR layer, consisting of a $3\times 3$ convolution, batch normalization (BN), and ReLU activation, is employed to extract spatial information. A DW convolutional layer is then integrated with residual connections to reinforce prior knowledge within the embedding process.
@@ 3.3.2 Temporal Information Branch
Most existing ViT-based approaches for extracting temporal dependencies via softmax attention often incur excessive computational complexity. While linear attention reduces computational overhead, it compromises model expressiveness, leading to inadequate representation capacity of the network. To this end, a novel attention mechanism, TAM, is proposed, integrating the advantages of both paradigms, enabling effective temporal dependency extraction in cloud imagery sequences as shown in Fig. 4 (c).
Different from the original agent attention, the proposed method does not employ standard projection matrices to derive the query ($Q$) and key ($K$) in the attention mechanism. Instead, customized convolutional kernels are constructed to establish independent adaptive receptive fields for each pixel in cloud imagery, thereby modeling long-range dependencies. These processes can be formally expressed as:
@@
$X_{K_{ij}}=\sum_{l=-1}^{1}\sum_{g=-1}^{1}E_{2+l,2+g}^{k}X_{i+l,j+g}$ (11)
where $E^{q}$ and $E^{k}\in\mathbb{R}^{T\times C\times 3\times 3}$ denotes a learnable projection matrix that aggregates features from the $3\times 3$ neighborhood of adjacent pixels into $X_{i,j}\in\mathbb{R}^{T\times C\times N}$. While analogous to the linear projection in ViT, this approach generates adaptively learned receptive fields, enhancing the capability of the network to perceive multi-scale variations in cloud imagery through dynamic spatial adaptation.
Subsequently, an agent token, $X_{A}\in\mathbb{R}^{T\times C\times n},n\ll N$, is generated via pooling operations. Notably, the tokens $n$ are set as a small hyper-parameter to achieve a linear computation complexity while maintaining global context modeling capability. Specifically, $X_{A}$ is first treated as the $Q$ to compute attention scores with $X_{K}$ and $X_{V}$, yielding the agent feature $A_{F}$. Then, $A_{F}$ is utilized as the $V$, while $X_{A}$ serves as the $K$, to perform a second attention computation with $X_{Q}$, producing the final attention feature $A$. These processes are formalized as:
$X_{A}=Pooling\left(X_{Q}\right)$ (12)
$A_{F}=Softmax\left(X_{A},\left(X_{K}\right)^{T}\right)X_{V}$ (13)
$A=Softmax\left(X_{Q},\left(X_{A}\right)^{T}\right)A_{F}$ (14)
where $Softmax$ denotes softmax attention function. In this way, we avoid expensive computational costs while preserving the information interaction between $Q$ and $K$. The feature $A$ is then processed through a DW convolutional layer combined with residual connections to obtain diversity in temporal feature information. Finally, following a DW convolution and BN, the output $X_{T}$ of the temporal branch is obtained. This process is expressed as:
$X_{T}=CB\left(Cat\left(DW\left(X_{V},A\right)\right)\right)$ (15)
@@
3.3.3 Dynamic Spatiotemporal Module
To ensure robust spatiotemporal dependency modeling in cloud image sequence extrapolation, we design a dynamic temporal module with TGM that applies weighted guidance from temporal flow information to spatial feature maps. Temporal dynamics enhance the capacity of the network to capture multi-scale contextual features and model long-term dependency.
[…]
@@ 3.4 Dynamic Update Decoding
The “ghosting effect” is a prevalent challenge in ground-based remote sensing cloud image sequence extrapolation tasks. The decoder of most existing methods is designed to fuse the same scale feature maps via a lateral connection. However, these methods lack temporal information guidance between hierarchical features and ignore the global feature correlation between the local information of different layers.
To this end, we propose a decoder with a DUM to enhance the ability of the method to capture global temporal flow correlations across different layers, which obtains the relationship of different level features with long-range dependencies. Specifically, we build a gate unit with initial temporal information $X_{T_{0}}$ to prevent the temporally guided spatiotemporal context $X_{D}$, derived from the USTM, from being diluted. As shown in Fig. 6, our TGM samples $X_{D}$ and $X_{T_{0}}$ for gate operation. One branch adopts the $X_{T_{0}}$ as the attention operator of the attention mechanism. The other adjusts the channel information adaptively by re-weighting the $X_{D}$ in the decoder through initial temporal flow information. Then, the features of the two branches are fused by a pointwise (PW) convolution to refine the fused information. The global features $D_{4}$ with temporal flow information are generated through the dynamic update decoding, which is conducive to obtaining the long-range dependence of image sequences. The $D_{4}$ will restore the feature scale as the third layer by passing the upsampling operation to obtain the third feature map. The similar processing of the next layer in the decoder will be repeated, and we can obtain the fused feature $D_{i}$. The entire process can be mathematically formulated as follows:
@@
$\mathrm{DUM}=Cat\left(X_{D},\mathrm{PWconv}\left(\mathrm{gate}\left({X_{T}}_{0},X_{D}\right)\right)\right)$ (17)
$\mathrm{gate}\left(X_{D},X_{T_{0}}\right)=\left(\omega_{1}\cdot X_{D}+b\right)\otimes\sigma\left(\omega_{2}\cdot X_{T_{0}}+c\right)$ (18)
where $\mathrm{PWconv}$ denotes the point-wise convolution, which is used to aggregate the features, $\mathrm{gate}$ denotes the gate unit. $\omega$ denotes the learning matrix, $b,c$ denotes the bias term, $\otimes$ denotes element-wise production, and $\sigma$ denotes the sigmod function. The $\mathrm{UP}$ operation is composed of two convolutional operations (the kernel size is 3) and bilinear interpolation with the scale factor is set to 2.

### v2

@@ 3.1 Formulation and architectural overview
The ground-based cloud image sequence extrapolation task operates on an input sequence of $T$ historical frames, denoted as $X_{t}^{T}=\left\{x_{i}\right\}_{t+1}^{T}$, where each frame $x_{i}\in\mathbb{R}^{C\times H\times W}$. The objective is to predict a subsequent sequence of $\tau$ future frames after $T$, $Y_{T+1}^{T+\tau}=\left\{y_{i}\right\}_{T+1}^{T+\tau}$, where $y_{i}\in\mathbb{R}^{C\times H\times W}$. Here, $C$, $H$, and $W$ represent the channel, height, and width dimensions, respectively. The model, parameterized by $\theta$, approximates a mapping function $F_{\theta}:X_{t}^{T}\rightarrow Y_{T+1}^{T+\tau}$. This mapping is optimized by maximizing the log-likelihood of the predicted cloud frames relative to their ground-truth counterparts. Formally, the optimization objective is defined as:
$\theta_{T}=arg\underset{\theta}{max}{\textstyle\sum_{i=T+1}^{T+\tau}}logP\left(y_{i}\mid x_{i};\theta\right).$ (1)
where $\theta_{T}$ represents the learnable parameters that align the predicted distribution with the physical reality of cloud formation dynamics.
@@ 3.2 Overview of structure
The proposed USF-Net adopts a sophisticated encoder-decoder configuration (Fig. 2a). The procedure of our USF-Net is described in Algorithm 1. To mitigate the inaccuracies arising from the scale-variant nature of cloud motion and the ghosting effects prevalent in existing decoders , USF-Net integrates a USTM at the bottleneck.
In the encoder, as shown in Fig. 2 (a) and (b), the input tensor $X_{t}^{T}\in\mathbb{R}^{B\times T\times C\times H\times W}$ traverses three Basic Layers. Each layer incorporates a $3\times 3$ Depth-Wise (DW) convolution for local feature extraction, followed by layer normalization, a secondary DW convolution, and a Squeeze-and-Excitation (SE) block to recalibrate channel dependencies. Residual connections facilitate gradient propagation, enriching the representation capabilities. We denote the feature map of the $i$-th layer as $f_{i}(i\in[1,2,3])$, with dimensions scaling to $T\times 2^{i-1}\cdot 64\cdot(H\times W)/2^{i+1}$. This process culminates in an intermediate feature map $X_{B}\in\mathbb{R}^{T\cdot C_{3}\cdot(H\times W)/16}$.
The USTM processes $X_{B}$ through dual pathways: a spatial branch that elicits multi-scale spatial features and a temporal branch that captures sequential dependencies. A DSM subsequently fuses these streams, leveraging temporal cues to guide spatial feature reconstruction.
In the decoder, we introduce the DUM to resolve the loss of contextual information often observed during upsampling. The DUM utilizes a gating mechanism to reweight initial temporal features, thereby refining the multi-scale spatiotemporal representations and preserving semantic coherence throughout the extrapolation horizon. Consequently, our architecture can be interpreted as a spatiotemporally unified framework optimized for ground-based cloud image extrapolation, balancing computational efficiency with prediction accuracy. The details of our method are as follows.
3.3 Unified spatioemporal module (USTM)
Accurate cloud extrapolation necessitates the simultaneous resolution of multi-scale spatial variations and non-stationary temporal dynamics. Conventional large-scale kernels often lack the adaptability required to capture fine-grained cloud textures, while standard self-attention mechanisms incur prohibitive quadratic computational costs. The USTM addresses these limitations by synergizing a spatial information branch (SiB), a temporal information branch (TiB), and a DSM (Fig. 2d).
@@ 3.3 Unified spatioemporal module (USTM)
3.3.1 Spatial information branch (SiB)
To accommodate the nonlinear and scale-variant nature of cloud formations, the SiB employs a SSM, as shown in Fig. 3. Inspired by the Large Kernel Selection (LSK) mechanism [LSK], the SSM dynamically adjusts the receptive field to align with the varying scales of cloud structures.
The SSM initially processes the input via a DW convolutional layer and a GeLU activation function, balancing efficiency with nonlinear expressiveness. We employ multi-scale convolutional kernels to adaptively select relevant spatial contexts. Let $K_{i}$ denote the $i$-th kernel; the resulting feature map $X_{K_{i}}$ is derived as:
$X_{K_{i}}=\mathrm{DW}\left(K_{i}\right).$ (2)
We utilize a decomposition strategy for large kernels, setting $K_{1}=3$ (dilation $d_{1}=1$) and $K_{2}=7$ (dilation $d_{2}=3$). The effective $\mathrm{RF}$ is expanded via:
$\mathrm{RF}_{1}$ $=k_{1}$ (3)
$\mathrm{RF}_{2}$ $=d_{2}\left(k_{2}-1\right)+\mathrm{RF}_{1}$ (4)
@@ 3.3.1 Spatial information branch (SiB)
This explicit decomposition mimics a large kernel ($K=21$) while minimizing parameter overhead. The multi-scale features are concatenated to form $X_{K}$, which encapsulates diverse receptive fields. To facilitate cross-spatial feature interaction, we apply channel-wise average and max pooling, followed by a convolution to generate the spatial interaction attention map $\widetilde{SA}$:
$X_{K}$ $=\mathrm{Cat}\left(X_{K_{1}},X_{K_{2}}\right)$ (5)
$\widetilde{SA}$ $=\mathrm{Conv}\left(\mathrm{Avg}\left(X_{K}\right),\mathrm{Max}\left(X_{K}\right)\right).$ (6)
@@
where $\mathrm{Cat}(\cdot)$ denotes the concatenation of the channel, $\mathrm{Avg}(\cdot)$ and $\mathrm{Max}(\cdot)$ denote the average pooling and max pooling, respectively.
Attention weights $\widehat{SA}_{i}$ for each scale are computed via a sigmoid function ($\sigma$). These weights modulate the decomposed feature maps, which are subsequently fused and convolved to produce the multi-scale spatial attention map $X_{SA}$. The final output $X_{S}$ is the element-wise product of the input $X_{B}$ and $X_{SA}$:
@@
$X_{SA}$ $=\mathrm{Conv}\left(\mathrm{Concat}\left(\widehat{SA}_{1}\ast X_{K_{1}},\widehat{SA}_{2}\ast X_{K_{2}}\right)\right)$ (8)
$X_{S}$ $=X_{SA}\ast X_{B}.$ (9)
where $\ast$ denotes the element-wise multiplication. By dynamically adjusting the receptive fields of spatial targets within the spatial branch, the proposed method effectively captures contextual information across varying cloud scales.
3.3.2 Temporal information branch (TiB)
Capturing the non-stationary motion of cloud sequences requires robust temporal modeling. Unlike recurrent networks, which suffer from sequential processing bottlenecks, or standard ViTs with quadratic complexity, we propose a TAM. This module integrates the high precision of Softmax attention with the linear complexity of Agent Attention, optimizing the trade-off between computational efficiency and representational capacity.
As shown in Fig. 4, The TiB comprises a Convolutional Embedding (CE) layer and stacked TAMs. Deviating from standard patch embedding, the CE layer utilizes a CNN to preserve spatial coherence. It consists of a CBR block ($3\times 3$ Conv, Batch Normalization, ReLU) followed by a DW convolution with residual connections.
@@ 3.3.2 Temporal information branch (TiB)
The TAM establishes long-range dependencies without the computational overhead of standard self-attention. Instead of projection matrices, we construct customized convolutional kernels to define adaptive receptive fields for each pixel. The Query ($Q$) and Key ($K$) projections are formulated as:
@@
$X_{K_{ij}}$ $=\sum_{l=-1}^{1}\sum_{g=-1}^{1}E_{2+l,2+g}^{k}X_{i+l,j+g}.$ (11)
where $E^{q},E^{k}\in\mathbb{R}^{T\times C\times 3\times 3}$ are learnable matrices aggregating local neighborhoods. To achieve linear complexity, we generate an agent token $X_{A}\in\mathbb{R}^{T\times C\times n}$ ($n\ll N$) via pooling. The attention computation proceeds in two stages: first between the Agent ($Q$) and the global context ($K,V$) to generate agent features $A_{F}$, and second between the global Query ($X_{Q}$) and the Agent ($K$) to distribute the attention:
$X_{A}$ $=\mathrm{Pooling}\left(X_{Q}\right)$ (12)
$A_{F}$ $=\mathrm{Soft}\left(X_{A},\left(X_{K}\right)^{T}\right)X_{V}$ (13)
$A$ $=\mathrm{Soft}\left(X_{Q},\left(X_{A}\right)^{T}\right)A_{F}.$ (14)
where $\mathrm{Soft}(\cdot)$ denotes softmax attention function. The resulting feature $A$ undergoes DW convolution and normalization to yield the final temporal embedding $X_{T}$.
$X_{T}$ $=\mathrm{BN}(\mathrm{DW}(\left(\mathrm{Concat}\left(\mathrm{DW}\left(X_{V},A\right)\right)\right)))$ (15)
@@
3.3.3 Dynamic spatiotemporal module (DSM)
As shown in Fig. 5, the outputs of the spatial and temporal branches are fused to generate a combined feature map $X_{F}\in\mathbb{R}^{T\times C\times H\times W}$, which is then enhanced via a DW convolutional layer and residual connection for enriched representation. The temporal guidance module is subsequently applied to implement time flow information guidance. Specifically, $X_{F}$ is split into two components, $X^{C}_{F}\in\mathbb{R}^{T\times C\times HW}=\mathrm{Conv}\left(X_{F}\right)$ and $X^{P}_{F}\in\mathbb{R}^{T\times C\times S^{2}}=\mathrm{Conv}\left(Pool\left(X_{F}\right)\right)$. The $X^{P}_{F}$ undergoes adaptive average pooling to aggregate spatial information into $S$ regions. Then, $X^{C}_{F}$ and $X^{C}_{F}$ are divided into $J$ groups along to channels to obtain $X^{C,J}_{F}\in\mathbb{R}^{T\times J\times\frac{C}{J}\times HW}=Re\left(X_{F}^{C}\right)$ and $X^{P,J}_{F}\in\mathbb{R}^{T\times J\times\frac{C}{J}\times S^{2}}=Re\left(X_{F}^{C}\right)$, respectively, where $Re\left(\cdot\right)$ denotes a reshape operation. A cross-correlation matrix, $X^{R}_{F}\in\mathbb{R}^{T\times HW\times S^{2}}$, is computed through matrix multiplication between corresponding groups, capturing inter-region contextual relationships. The key idea is to represent inter-region contextual relationships via $J$ group vectors, enabling the learning of dynamic convolution kernels from $X^{R}_{F}$. Long-term dependencies are dynamically modulated by propagating contextual information across correlated regions. Subsequently, $J$ dynamic convolution kernels of size $K\times K$ are generated by mapping $X^{R}_{F}$ through a learnable linear layer $W\in\mathbb{R}^{S^{2}\times K^{2}}$, producing spatiotemporal tokens that encode regional context from the correlation matrix. The feature $X_{F}$ is also divided into $J$ channel groups, which are then convolved with the reshaped kernels to share spatiotemporal dependencies, yielding the dynamically modulated feature $X_{F^{\prime}}$. The output $X_{U}$ of the TGM is obtained by combining $X_{F}$ and $X_{F^{\prime}}$. Finally, $X_{U}$ is processed through a normalization layer and a convolutional layer to generate the output $X_{D}$ of the dynamic spatiotemporal module.
@@ 3.4 Dynamic update decoding
A pervasive challenge in cloud sequence extrapolation is the ghosting effect, characterized by blurred predictions due to information decay in the decoder. Standard decoders using simple lateral connections often fail to preserve high-frequency temporal details. To rectify this, we introduce a decoder equipped with DUM.
The DUM functions as a gated recurrent unit that reinforces the temporally guided context $X_{D}$ (from USTM) with the initial temporal information $X_{T_{0}}$ from the TAM, as shown in Fig. 6. This mechanism prevents the dilution of temporal cues during upsampling. Then, the features of the two branches are fused by a pointwise (PW) convolution to refine the fused information. The global features $D_{4}$ with temporal flow information are generated through the dynamic update decoding, which is conducive to obtaining the long-range dependence of image sequences. The $D_{4}$ will restore the feature scale as the third layer by passing the upsampling operation to obtain the third feature map. The similar processing of the next layer in the decoder will be repeated, and we can obtain the fused feature $D_{i}$. The entire process can be mathematically formulated as follows:
@@
$\mathrm{DUM}$ $=\mathrm{Concat}\left(\mathrm{PW}\left(X_{D},\mathrm{gate}\left({X_{T}}_{0},X_{D}\right)\right)\right)$ (17)
$\mathrm{gate}\left(X_{D},X_{T_{0}}\right)$ $=\left(\omega_{1}\cdot X_{D}+b\right)\ast\sigma\left(\omega_{2}\cdot X_{T_{0}}+c\right).$ (18)
where $\mathrm{PW}(\cdot)$ denotes the point-wise convolution, which is used to aggregate the features, $\mathrm{gate}$ denotes the gate unit. $\omega$ denotes the learning matrix, $b,c$ denotes the bias term, and $\sigma$ denotes the sigmod function. The $\mathrm{UP}$ operation is composed of two convolutional operations (the kernel size is 3) and bilinear interpolation with the scale factor is set to 2.